OpenStudy (anonymous):

Solve the given equation. 6^3n+6= 1,296

OpenStudy (anonymous):

1,296=6^4... other steps are analog like in previous example

OpenStudy (anonymous):

okay thanks :) see thats what i have a problem in. i cant seem to find the quare roots of big numbers

OpenStudy (anonymous):

wait what do i do with the 6^3?

OpenStudy (anonymous):

6^(3n+3)=6^4 so 3n+3=4...

OpenStudy (anonymous):

ok. so i dont put 216=6^3

OpenStudy (anonymous):

why would you

OpenStudy (anonymous):

idk... because 6^3=216

OpenStudy (anonymous):

so what? you want such base with you can solve exponential equation... so on left side you need same base like on right side...

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

oh wait one last question lol.. how did you get the +3 in 6^(3n+3)?